# Jailbreak Attack Initializations as Extractors of Compliance Directions

### v2

@@ C.1.4 Attack Evaluation
In Figure 24, we compare the $ASR$ of our initializations vs. standard and random using a GPT-4 judge for the $ASR$-Recheck calculation on the output, according to Liu et al. (2023). This is relevant because the current evaluation of $GCG$ is outdated, results and further details represented in C.5.3.
@@ C.3 CRI Optimization - HarmBench
Table 2: Attack-success rate ($ASR$%) for each model, initialization, and cluster.
@@ C.4.1 Full Experiments Tables - HarmBench and AdvBench
We evaluate the open-source LLMs as mentioned in Section C.1.2.
@@
Table 3: Universal Attack Trained on Vicuna(Bottom) and Llama-2(Top). Comparison of $25\text{-}CRI$ and $1\text{-}CRI$ to Standard and Random initialization on $GCG\text{-}M$ transfer attacks over the $AdvBench$ dataset. The source model is Vicuna, and $ASR$ (%) is presented for each target model.
@@
In Figure 9, we present evaluations on the $GCG$ attack using two younger models: DeepSeek-LLM-7B-Chat and Falcon3-7B-Instruct. The DeepSeek models seems to be the least robust out of the four examined, demonstrating $100$ $ASR$ quickly using all initializations, with ours taking the lead. When examining Falcon, we can notice a clear difference between the uninformative initializations and ours, where all of ours achieve higher $ASR$ way more quickly.
@@
Figure 9: Comparison of $K\text{-}CRI$ ($K=1,5,25$) to standard and random initialization on the $GCG$ attack over the $HarmBench$ dataset. The attacks’ $ASR$ are presented on Falcon3-7B-Instruct (left) and deepseek-llm-7b-chat (right).
@@
Figure 10: Extra $ASR$ vs. steps results on the embedding attack with three more models on the $HarmBench$ dataset.
@@
In Figure 11, we present evaluations on Llama-2, Vicuna and Llama-3. Our initialization prove superior in both Llama models comparative to the uninformative initializations. On Vicuna, which is a less robust model, all initializations converge quickly to $100$ $ASR$. But, our initializations still converge way more quickly, demonstrating its advantage even here, in a less robust setting. In Figures 12 and 13, we present evaluations on the same models on the $AutoDAN$ attack variants ($GA,HGA$), when using $1\text{-}CRI$ vs. its standard initialization. On Llama-2, we can notice an improvement in the $ASR$ in both variants, converging more quickly to a higher $ASR$. On Llama-3, which is surprisingly less robust than Llama-2 in the attack, we can notice a slight faster convergence to $100$ $ASR$. On Vicuna, the attacks converge to $100$ $ASR$ very quickly, noticing little to no difference between our initialization and the standard one.
@@
Figure 11: Comparison of $K\text{-}CRI$ ($K=1,5,25$) to standard and random initialization on the $GCG$ attack over the $AdvBench$ dataset. The attacks’ $ASR$ are presented on Llama-2 (left), Vicuna (middle), and Llama-3 (right).
@@
Figure 12: Comparison of $25\text{-}CRI$ to standard initialization on the $AutoDAN\text{-}GA$ attack over the $AdvBench$ dataset. The attacks’ $ASR$ are presented on Llama-2 (left), Vicuna (middle), and Llama-3 (right).
@@
Figure 13: Comparison of $25\text{-}CRI$ to standard initialization on the $AutoDAN\text{-}HGA$ attack over the $AdvBench$ dataset. The attacks’ $ASR$ are presented on Llama-2 (left), Vicuna (middle), and Llama-3 (right).
@@
Figure 14: Embedding attack results on $AdvBench$. We compare $K\text{-}CRI$ ($K{=}1,5,25$) to standard and random initialization across five models. Top row (left to right): Llama-2, Vicuna, Llama-3. Bottom row (left to right): Llama-3.1, Mistral-7B.
@@
Figure 15: $LFS$ versus steps-to-success for four initialization strategies on Vicuna in the $GCG$ attack. The dashed line is the least-squares regression fit across all points (Pearson $r=0.31$, $p=1\times 10^{-5}$).
@@
Figure 16: $LFS$ versus the number of optimization steps required for the embedding attack across five LLMs. Llama-2 ($r=0.62,\;p=4\times 10^{-17}$); Vicuna ($r=0.70,\;p=2\times 10^{-23}$); Llama-3-8B ($r=0.32,\;p=7\times 10^{-5}$); Llama-3.1-8B ($r=0.39,\;p=10^{-6}$); Mistral-7B ($r=0.51,\;p=4\times 10^{-11}$).
@@ C.5.1 Batch Size
In this part, we experiment with the batch size hyperparameter of the $GCG$ attack, which governs computational resources per attack iteration. A larger batch size allows considering more potential token replacements at each iteration. In Figure 17, we evaluate on the $HarmBench$ dataset, and in Figure 18, on $AdvBench$. We observe that for both datasets, using a small batch size of $16$, the $1\text{-}CRI$ initialization underperforms compared to other $CRI$s and the baseline. However, performance significantly improves, becoming comparable or superior to other initializations, when the batch size increases. Additionally, the baseline initialization severely underperforms relative to $25\text{-}CRI$ and $5\text{-}CRI$ at small batch sizes ($16$ and $32$), and only becomes somewhat comparable, yet still inferior, at larger batch sizes ($64$). This highlights the significance of our framework under varying computational resources.
@@
Figure 17: Comparison of $K\text{-}CRI$ ($K=1,5,25$) to standard and random initialization on the $GCG$ attack over the $HarmBench$ dataset. The attacks’ $ASR$ are presented on Llama-2 (left) and Vicuna (right). Across different batch sizes: 16 (top), 32 (center), and 64 (bottom).
@@
Figure 18: Comparison of $K\text{-}CRI$ ($K=1,5,25$) to standard and random initialization on the $GCG$ attack over the $AdvBench$ dataset. The attacks’ $ASR$ are presented on Llama-2 (left) and Vicuna (right). Across different batch sizes: 16 (top), 32 (center), and 64 (bottom).
@@ C.5.2 CRI Cost
$C_{\text{total}}=C_{CRI}+N_{test}\cdot C_{CRI\text{-deploy}}\ll N_{test}\cdot C_{\text{base}}\quad\text{when }N_{test}\gg K.$
Ablation study of the number of steps used to create the initialization on $GCG$ in the $HarmBench$ dataset can be found in Figures 19, 20 and 21. We notice that quite consistently, the more steps used for training the $CRI$ set, the better the results in deployment are, with not a large difference between $400$ and $500$ initialization steps.
@@
To evaluate the generalization capacity of $CRI$, we train the initialization set $\mathcal{T}_{K\text{-}CRI}$ on $S_{FT}^{\text{AdvBench}}$ and directly deploy it on $S_{FT}^{\text{HarmBench}}$ without retraining. Experiments on both Llama-2 and Vicuna show that $CRI$ maintains high $ASR$ across datasets, suggesting that the compliance-refusal structure captured by $CRI$ is model- and dataset-agnostic. This transferability further reduces the need for repeated initialization training, offering near-zero overhead in cross-dataset settings. In Figure 22, we compare using the same dataset for initialization and testing vs. using separate datasets. Here we test on the $HarmBench$ dataset, and compare using it for initialization vs. using $AdvBench$ for initialization.
@@
Figure 22: Attack success rates on $HarmBench$: in-dataset initialization versus cross-dataset initialization from $AdvBench$. Llama-2 (left) and Vicuna (right).
@@
To better evaluate $CRI$’s effectiveness under fair cost conditions, we compare it to baseline attacks executed with a proportionally increased computational budget. Specifically, we benchmark $CRI$ against baselines that are allowed twice the number of optimization iterations (e.g., 500 steps). In Figure 23, we compared $25-CRI$ with baselines over a proportionally increased computational budget.
Figure 23: Comparison of $25-CRI$ and running the standard attack for an extended amount of steps.
@@ C.5.3 Additional Studies, Attacks Transfer, Defenses, Evaluations And Integration with other improvements
We evaluated the transferability of initialization by extracting the 1-CRI initialization set from the GCG individual attack Liu et al. (2022) and using it to initialize the Embedding attack Schwinn et al. (2024), we choose the model $deepseek-llm-7b-chat$ as we find it new and robust for jailbreak attacks. While partial improvements were observed, the effectiveness was limited, suggesting that the optimization dynamics of these two distinct attack paradigms differ significantly. This highlights that while CRI can generalize across certain frameworks, its performance is sensitive to the underlying optimization mechanisms
@@
In our work, we adhere to the evaluation setting used in the original attack baseline to ensure coherent results, however, we acknowledge that more sophisticated metrics could further enhance robustness. We have extended our evaluation as recommended to include a GPT-4-based judgment setting Liu et al. (2024a). Following the protocol from Liu et al. (2024a), GPT-4 is used to assess the $ASR$ of generated jailbreak prompts. Our results show that $CRI$ continues to improve $ASR$ in this setup. Furthermore, the best-performing initialization remains consistent across both the keyword-matching and GPT-4-based evaluation frameworks, reinforcing the reliability and transferability of $CRI$-enhanced prompts, the results are demonstrated in Figure 24.
@@
We present additional evaluations of our approach in more challenging settings. Specifically, we integrate BOOST Yu et al. into our attack framework, using a pre-trained initialization set in place of the standard $GCG$ initialization. The resulting variant—combining both $CRI$ and BOOST—achieves the highest $ASR$, demonstrating that $CRI$ is compatible with BOOST without requiring retraining. To compare against baselines in the setting proposed by the original BOOST authors, we use a keyword-matching evaluation method. In Figure 25, we compare our initializations to using BOOST Yu et al. .
@@
Warning: The following sections contains potentially offensive and harmful text.
@@ Appendix F CRI Set Example
Figure 29: $CRI$ set for Llama-2 on the $GCG$ attack.
e

### v3

@@ C.1.4 Attack Evaluation
In Figure˜26, we compare the $ASR$ of our initializations vs. standard and random using a GPT-4 judge for the ASRASR-Recheck calculation on the output, according to Liu et al. (2023). This is relevant because the current evaluation of GCGGCG is outdated, results and further details represented in C.5.3.
@@ C.3 CRI Optimization - HarmBench
Table 2: Attack-success rate (ASRASR%) for each model, initialization, and cluster.
@@ C.4.1 Full Experiments Tables - HarmBench and AdvBench
We evaluate the open-source LLMs as mentioned in Section˜C.1.2.
@@
Table 3: Universal Attack Trained on Vicuna(Bottom) and Llama-2(Top). Comparison of $25\text{-}CRI$ and $1\text{-}CRI$ to Standard and Random initialization on $GCG\text{-}M$ transfer attacks over the $AdvBench$ dataset. The source model is Vicuna, and ASRASR (%) is presented for each target model.
@@
In Figure˜10, we present evaluations on the $GCG$ attack using two younger models: DeepSeek-LLM-7B-Chat and Falcon3-7B-Instruct. The DeepSeek models seems to be the least robust out of the four examined, demonstrating $100$ ASRASR quickly using all initializations, with ours taking the lead. When examining Falcon, we can notice a clear difference between the uninformative initializations and ours, where all of ours achieve higher $ASR$ way more quickly.
@@
Figure 10: Comparison of $K\text{-}CRI$ ($K=1,5,25$) to standard and random initialization on the $GCG$ attack over the $HarmBench$ dataset. The attacks’ ASRASR are presented on Falcon3-7B-Instruct (left) and deepseek-llm-7b-chat (right).
@@
Figure 11: Extra ASRASR vs. steps results on the embedding attack with three more models on the $HarmBench$ dataset.
@@
In Figure˜12, we present evaluations on Llama-2, Vicuna and Llama-3. Our initialization prove superior in both Llama models comparative to the uninformative initializations. On Vicuna, which is a less robust model, all initializations converge quickly to $100$ ASRASR. But, our initializations still converge way more quickly, demonstrating its advantage even here, in a less robust setting. In Figures˜13 and 14, we present evaluations on the same models on the $AutoDAN$ attack variants ($GA,HGA$), when using $1\text{-}CRI$ vs. its standard initialization. On Llama-2, we can notice an improvement in the $ASR$ in both variants, converging more quickly to a higher $ASR$. On Llama-3, which is surprisingly less robust than Llama-2 in the attack, we can notice a slight faster convergence to $100$ ASRASR. On Vicuna, the attacks converge to $100$ ASRASR very quickly, noticing little to no difference between our initialization and the standard one.
@@
Figure 12: Comparison of $K\text{-}CRI$ ($K=1,5,25$) to standard and random initialization on the $GCG$ attack over the $AdvBench$ dataset. The attacks’ ASRASR are presented on Llama-2 (left), Vicuna (middle), and Llama-3 (right).
@@
Figure 13: Comparison of $25\text{-}CRI$ to standard initialization on the $AutoDAN\text{-}GA$ attack over the $AdvBench$ dataset. The attacks’ ASRASR are presented on Llama-2 (left), Vicuna (middle), and Llama-3 (right).
@@
Figure 14: Comparison of $25\text{-}CRI$ to standard initialization on the $AutoDAN\text{-}HGA$ attack over the $AdvBench$ dataset. The attacks’ ASRASR are presented on Llama-2 (left), Vicuna (middle), and Llama-3 (right).
@@
Figure 15: Embedding attack results on $AdvBench$. We compare $K\text{-}CRI$ (K=1,5,25K{=}1,5,25) to standard and random initialization across five models. Top row (left to right): Llama-2, Vicuna, Llama-3. Bottom row (left to right): Llama-3.1, Mistral-7B.
@@
Figure 16: Distribution of per-prompt Spearman correlation of LFSLFS and number of steps-to-success, on Vicuna. (Median $r=0.46$, $p=5.73\times 10^{-2}$).
Figure 17: LFSLFS versus steps-to-success for four initialization strategies on Vicuna in the GCGGCG attack. The dashed line is the least-squares regression fit across all points (Pearson $r=0.31$, $p=1\times 10^{-5}$).
@@
Figure 18: LFSLFS versus the number of optimization steps required for the embedding attack across five LLMs, Pearson correlation. Llama-2 ($r=0.62,\;p=4\times 10^{-17}$); Vicuna ($r=0.70,\;p=2\times 10^{-23}$); Llama-3-8B ($r=0.32,\;p=7\times 10^{-5}$); Llama-3.1-8B ($r=0.39,\;p=10^{-6}$); Mistral-7B ($r=0.51,\;p=4\times 10^{-11}$).
@@ C.5.1 Batch Size
In this part, we experiment with the batch size hyperparameter of the $GCG$ attack, which governs computational resources per attack iteration. A larger batch size allows considering more potential token replacements at each iteration. In Figure˜19, we evaluate on the $HarmBench$ dataset, and in Figure˜20, on $AdvBench$. We observe that for both datasets, using a small batch size of $16$, the $1\text{-}CRI$ initialization underperforms compared to other $CRI$s and the baseline. However, performance significantly improves, becoming comparable or superior to other initializations, when the batch size increases. Additionally, the baseline initialization severely underperforms relative to $25\text{-}CRI$ and $5\text{-}CRI$ at small batch sizes ($16$ and $32$), and only becomes somewhat comparable, yet still inferior, at larger batch sizes ($64$). This highlights the significance of our framework under varying computational resources.
@@
Figure 19: Comparison of $K\text{-}CRI$ ($K=1,5,25$) to standard and random initialization on the $GCG$ attack over the $HarmBench$ dataset. The attacks’ ASRASR are presented on Llama-2 (left) and Vicuna (right). Across different batch sizes: 16 (top), 32 (center), and 64 (bottom).
@@
Figure 20: Comparison of $K\text{-}CRI$ ($K=1,5,25$) to standard and random initialization on the $GCG$ attack over the $AdvBench$ dataset. The attacks’ ASRASR are presented on Llama-2 (left) and Vicuna (right). Across different batch sizes: 16 (top), 32 (center), and 64 (bottom).
@@ C.5.2 CRI Cost
$C_{\text{CRI-total}}=C_{CRI}+N_{test}\cdot C_{CRI\text{-deploy}}\ll C_{\text{total}}=N_{test}\cdot C_{\text{base}}\quad\text{when }N_{test}\gg K.$
Runtime Analysis
On our hardware (Intel(R) Xeon(R) CPU and NVIDIA L40S GPU), we find that the GCG attack steps take different GPU times on different models, as expected. On Llama-2-7B, a single iteration of GCG takes approximately $3\times 10^{-4}$ GPU hours. As such, our computational analysis is as follows (Table˜1):
$C_{CRI}\approx 500[step]\times 25[prompts]\times 3\times 10^{-4}[hours/step]=3.75[hours]$
where each initialization is trained for $500$ iterations, and there are $25$ prompts in the initialization set.
$C_{base}\approx 212.16[step]\times 3\times 10^{-4}[hours/step]=0.064[hours]$
$C_{CRI-deploy}\approx 22.64[step]\times 3\times 10^{-4}[hours/step]=0.0068[hours]$
where $212.16$ and $22.64$ are the average steps to success of the standard and $25\text{-}CRI$ initializations.
On a test set of $500$ prompts:
$C_{CRI-total}=3.75[hours]+500[prompts]\times 0.0068[hours]=7.15[hours]$
$C_{total}=500[prompts]\times 0.064[hours]=32[hours]$
That entails a savings of approximately 25 GPU hours, in addition to the attack success rate improvements.
Ablation
Ablation study of the number of steps used to create the initialization on $GCG$ in the $HarmBench$ dataset can be found in Figures˜21, 22 and 23. We notice that quite consistently, the more steps used for training the $CRI$ set, the better the results in deployment are, with not a large difference between $400$ and $500$ initialization steps.
@@
To evaluate the generalization capacity of $CRI$, we train the initialization set $\mathcal{T}_{K\text{-}CRI}$ on $S_{FT}^{\text{AdvBench}}$ and directly deploy it on $S_{FT}^{\text{HarmBench}}$ without retraining. Experiments on both Llama-2 and Vicuna show that $CRI$ maintains high $ASR$ across datasets, suggesting that the compliance-refusal structure captured by $CRI$ is model- and dataset-agnostic. This transferability further reduces the need for repeated initialization training, offering near-zero overhead in cross-dataset settings. In Figure˜24, we compare using the same dataset for initialization and testing vs. using separate datasets. Here we test on the $HarmBench$ dataset, and compare using it for initialization vs. using $AdvBench$ for initialization.
@@
Figure 24: Attack success rates on HarmBenchHarmBench: in‑dataset initialization versus cross‑dataset initialization from AdvBenchAdvBench. Llama‑2 (left) and Vicuna (right).
@@
To better evaluate $CRI$’s effectiveness under fair cost conditions, we compare it to baseline attacks executed with a proportionally increased computational budget. Specifically, we benchmark $CRI$ against baselines that are allowed twice the number of optimization iterations (e.g., 500 steps). In Figure˜25, we compared 25CRI25-CRI with baselines over a proportionally increased computational budget.
Figure 25: Comparison of 25CRI25-CRI and running the standard attack for an extended amount of steps.
@@ C.5.3 Additional Studies, Attacks Transfer, Defenses, Evaluations And Integration with other improvements
We evaluated the transferability of initialization by extracting the 1-CRI initialization set from the GCG individual attack Liu et al. (2022) and using it to initialize the Embedding attack Schwinn et al. (2024), we choose the model deepseekllm7bchatdeepseek-llm-7b-chat as we find it new and robust for jailbreak attacks. While partial improvements were observed, the effectiveness was limited, suggesting that the optimization dynamics of these two distinct attack paradigms differ significantly. This highlights that while CRI can generalize across certain frameworks, its performance is sensitive to the underlying optimization mechanisms
@@
In our work, we adhere to the evaluation setting used in the original attack baseline to ensure coherent results, however, we acknowledge that more sophisticated metrics could further enhance robustness. We have extended our evaluation as recommended to include a GPT-4-based judgment setting Liu et al. (2024a). Following the protocol from Liu et al. (2024a), GPT-4 is used to assess the $ASR$ of generated jailbreak prompts. Our results show that $CRI$ continues to improve $ASR$ in this setup. Furthermore, the best-performing initialization remains consistent across both the keyword-matching and GPT-4-based evaluation frameworks, reinforcing the reliability and transferability of $CRI$-enhanced prompts, the results are demonstrated in Figure˜26.
@@
We present additional evaluations of our approach in more challenging settings. Specifically, we integrate BOOST Yu et al. (2024a) into our attack framework, using a pre-trained initialization set in place of the standard $GCG$ initialization. The resulting variant—combining both $CRI$ and BOOST—achieves the highest $ASR$, demonstrating that $CRI$ is compatible with BOOST without requiring retraining. To compare against baselines in the setting proposed by the original BOOST authors, we use a keyword-matching evaluation method. In Figure˜27, we compare our initializations to using BOOST Yu et al. (2024a).
@@
Warning: The following sections contain potentially offensive and harmful text.
@@ Appendix F CRI Set Example
Figure 31: CRICRI set for Llama-2 on the GCGGCG attack.